OpenStudy (laddiusmaximus):

this is a curve sketching question. I dont quite know how to start it. y=x^2(x+1)(3-2x)

OpenStudy (anonymous):

Since it already looks factored, I would start finding the zeros and then marking them down on your graph paper, before we move on to finding critical values, 2nd dervs,

OpenStudy (laddiusmaximus):

ok well i got as far as (x+1)=0 that would make x=-1

OpenStudy (laddiusmaximus):

and 3-2x=0 would make x= 3/2 right?

OpenStudy (anonymous):

sorry, I got dc'ed. I think you missed a zero as well.

OpenStudy (anonymous):

next you would take the 1st derivative and then solve IT for zero (not the previous function). That will give you the locations where the tangent = 0. We still don't know if that is a min or max yet though. For that I think it is easiest to take the 2nd derivative and then plug those critical values (where tangent = 0) into it.